OpenStudy (anonymous):

Convert to a percent. 0.375% 3.75% 37.5%

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

oops 3/8

OpenStudy (anonymous):

with out a calculator we know that 1/4=.25 like a quarter of a dollar so 1/8 is half that amount to it is .125 so .125*3=? and then move the decimal point over two to the right and you have a percent.

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 }= \frac{ \frac{ 3 }{ 2 } }{ 4 }=\frac{ 25*\frac{ 3 }{ 2 } }{ 100 }=\frac{ 75 }{2 } percent =37.5\]